# ToxVidLLM: A Multimodal LLM-based Framework for Toxicity Detection in Code-Mixed Videos

### v1

@@ 6 Conclusion and Future Works
In an ever-evolving internet landscape, where videos have become the predominant form of content, the challenge of detecting toxic content, especially in low-resource code-mixed languages, is more critical than ever. We introduce ToxCMM, a pioneering benchmark dataset featuring code-mixed videos for toxic content detection. Our proposed LLM-based advanced multimodal framework (ToxVidLLM) achieved remarkable results, emphasizing the significance of combining text, audio, and video modalities. It is worth noting that, among the individual modalities, transformer encodings of text prove to be particularly effective in detecting toxic videos. Beyond toxicity, the ToxCMM dataset includes two additional labels, sentiment, and severity, offering a comprehensive resource for further exploration in sentiment analysis within low-resource code-mixed videos. Our research emphasizes AI’s role in fostering a respectful online environment and promoting civility against toxic speech in video data.
@@ 7 Limitations
3. The proposed ToxVidLLM model fine-tunes two encoder modules, an LLM, and additional modules, requiring a considerable amount of GPU memory for training. Due to computational limitations, we were unable to experiment even with parameter-efficient fine-tuning methods (PEFT) like LoRA Hu et al. (2021) or Quantized-LoRA Dettmers et al. (2023) for models containing billions of parameters like OpenHathi-7B [^7] https://huggingface.co/sarvamai/OpenHathi-7B-Hi-v0.1-Base, Airavata-7B Gala et al. (2024), Llama 2-7B Touvron et al. (2023), or Mistral-7B Jiang et al. (2023). However, since our model is versatile, those with ample GPU resources can easily substitute larger models into the textual side, potentially achieving enhanced performance specifically for video classification tasks.
@@ Appendix A Qualitative Analysis
Figure 4: Human annotation Vs. model’s prediction for qualitative analysis with different settings
@@
(i) In the first sample (a), the textual modality ("Let us tell you our status now") lacks offensive words but constitutes a toxic utterance with negative sentiment and a severity score of 1 based on acoustic and visual expression. The user is threatening someone. Both the best baseline model and our proposed model with two modalities settings (ToxVidLLM2M) make incorrect predictions across all tasks. However, considering three modalities variants (ToxVidLLM3M) allows accurate prediction of all classes. This observation underscores our proposed model’s enhanced comprehension of diverse modalities, demonstrating that incorporating audio and visual cues with text provides a superior understanding of video data.
@@
(iii) In the third example, both baseline and single-task variants of the proposed model (ToxVidLLMST) inaccurately predict all labels, while the multitask model (ToxVidLLMMT) correctly identifies all classes. This example is toxic due to offensive language and derogatory assumptions about someone’s grandfather. The terms "sidhi lagti hogi" and "khud hi chadh jate hai" are disrespectful, implying negative sentiment. Task-specific layers in the multitask framework aid in correctly identifying true sentiment, leading to the accurate identification of toxicity and severity labels in the given post.

### v2

@@ 6 Conclusion and Future Works
In an ever-evolving internet landscape, where videos have become the predominant form of content, the challenge of detecting toxic content, especially in low-resource code-mixed languages, is more critical than ever. We introduce ToxCMM, a pioneering benchmark dataset featuring code-mixed videos for toxic content detection. Our proposed LM-based advanced multimodal framework (ToxVidLM) achieved remarkable results, emphasizing the significance of combining text, audio, and video modalities. It is worth noting that, among the individual modalities, transformer encodings of text prove to be particularly effective in detecting toxic videos. Beyond toxicity, the ToxCMM dataset includes two additional labels, sentiment, and severity, offering a comprehensive resource for further exploration in sentiment analysis within low-resource code-mixed videos. Our research emphasizes AI’s role in fostering a respectful online environment and promoting civility against toxic speech in video data.
@@ 7 Limitations
3. The proposed ToxVidLM model fine-tunes two encoder modules, an LM, and additional modules, requiring a considerable amount of GPU memory for training. Due to computational limitations, we were unable to experiment even with parameter-efficient fine-tuning methods (PEFT) like LoRA Hu et al. (2021) or Quantized-LoRA Dettmers et al. (2023) for models containing billions of parameters like OpenHathi-7B [^7] https://huggingface.co/sarvamai/OpenHathi-7B-Hi-v0.1-Base, Airavata-7B Gala et al. (2024), Llama 2-7B Touvron et al. (2023), or Mistral-7B Jiang et al. (2023). However, since our model is versatile, those with ample GPU resources can easily substitute larger models into the textual side, potentially achieving enhanced performance specifically for video classification tasks.
@@ Appendix A Qualitative Analysis
Figure 4: Human annotation versus model’s prediction for qualitative analysis with different settings
@@
(i) In the first sample (a), the textual modality ("Let us tell you our status now") lacks offensive words but constitutes a toxic utterance with negative sentiment and a severity score of 1 based on acoustic and visual expression. The user is threatening someone. Both the best baseline model and our proposed model with two modalities settings (ToxVidLM2M) make incorrect predictions across all tasks. However, considering three modalities variants (ToxVidLM3M) allows accurate prediction of all classes. This observation underscores our proposed model’s enhanced comprehension of diverse modalities, demonstrating that incorporating audio and visual cues with text provides a superior understanding of video data.
@@
(iii) In the third example, both baseline and single-task variants of the proposed model (ToxVidLMST) inaccurately predict all labels, while the multitask model (ToxVidLMMT) correctly identifies all classes. This example is toxic due to offensive language and derogatory assumptions about someone’s grandfather. The terms "sidhi lagti hogi" and "khud hi chadh jate hai" are disrespectful, implying negative sentiment. Task-specific layers in the multitask framework aid in correctly identifying true sentiment, leading to the accurate identification of toxicity and severity labels in the given post.